OpenStudy (anonymous):

I know this is very simple but I haven't done these for some times. 5x=0 is the question

OpenStudy (anonymous):

you are solving for x? Isolate X on the side by itself. Easiest way - divide both sides by 5

OpenStudy (anonymous):

its a quadratic equation. Yes x

OpenStudy (anonymous):

would it be x=0?

OpenStudy (texaschic101):

5x = 0 (divide by 5) 5x/5 = 0/5 x = 0 you are correct :)

OpenStudy (anonymous):

yeah! I still got it! Thx man

OpenStudy (anonymous):

What you typed is not quadratic, that is an linear equation

OpenStudy (texaschic101):

you still got it jaredfleagle :)

OpenStudy (anonymous):

so then if another question would be 3x(x+12)=0. The Answer would be x=-12 or x=0?

OpenStudy (texaschic101):

3x(x + 12) = 0 x = 0 x + 12 = 0

OpenStudy (anonymous):

alright thx. I don't have it anymore haha. Thx again